### 10 Apr 2024 v3

# Nonlocal Contributions to the Turbulent Cascade in Magnetohydrodynamic Plasmas

Abstract
We present evidence for nonlocal contributions to the turbulent energy cascade in magnetohydrodynamic (MHD) plasmas. Therefore, we revisit a well-known result derived directly from the MHD equations, i.e., the Politano & Pouquet (P&P) law for the transfer of kinetic and magnetic energy in scale. We propose adding a term that accounts for nonlocal transfer and represents the influence of fluctuations from large scales due to the Alfvén effect. Supported by direct numerical simulations of homogeneous and isotropic MHD turbulence, we verify that in some plasma configurations, neglecting the additional nonlocal term might consistently overestimate energy dissipation rates and, thus, the contributions of turbulent energy dissipation potentially affecting solar wind heating; a central puzzle in space plasma physics that motivates the present work.
I Introduction
The physical mechanisms underlying the solar wind - a continuous flow of charged particles emitted from the solar corona - have been hypothesized since the beginning of the past century. Based on Biermann’s observations of comet tail motion relative to the Sun Biermann (1951), Parker provided the first comprehensive theoretical description of the solar wind as a rapidly expanding outer coronal atmosphere overcoming the Sun’s gravitational field Parker (1958). Assuming an adiabatic solar wind expansion, Parker also considered the radial temperature profile $T(r)$ as a function of the heliocentric distance $r$. Nonetheless, a purely adiabatic expansion of the solar wind, which suggests a radial temperature profile of proton temperature as $T(r)\sim r^{-4/3}$, differs from actual solar wind plasma measurements that exhibit a much slower decay with increasing distances $r$ Matthaeus et al. (1999); Richardson and Smith (2003); Hundhausen (2012).
A possible mechanism that would supply heat to the expanding solar wind - and thus explain the slower decay of the temperature profile - is the dissipation of turbulent kinetic and magnetic energy into heat predicted by a phenomenological description of magnetohydrodynamic (MHD) turbulence Marino et al. (2008) (we also refer the reader to the recent review article on scaling laws in solar wind turbulence Marino and Sorriso-Valvo (2023)). The concept of turbulent cascades, which has been put forth in the context of hydrodynamic turbulence by Kolmogorov Kolmogorov (1941), Heisenberg Heisenberg (1948), von Weizsäcker v. Weizsäcker (1948), and Onsager Onsager (1949), highlights the fact that turbulent motions are essentially transport processes of energy in scale. The existence of energy cascades in fluid and plasma turbulence has been assessed experimentally (e.g., in channel flows and fusion plasmas Danaila et al. (2001); Conway (2008)), as well as by atmospheric Vierinen et al. (2019), oceanic Balwada et al. (2022), and space plasma measurements Telloni (2023). A central quantity in phenomenological descriptions of such cascade processes is the averaged energy dissipation rate $\langle\varepsilon_{kin}\rangle$, which is assumed to entirely characterize the transfer of energy from the injection scale, where the turbulent flow is stirred, to small scales where energy is dissipated into heat.
In a conducting fluid, the rate at which kinetic and magnetic energy is dissipated is determined by the local kinetic and magnetic energy dissipation rates
$\varepsilon_{kin}({\bf x},t)=$ $\frac{\nu}{2}\sum_{i,k}\left(\frac{\partial u_{i}({\bf x},t)}{\partial x_{k}}+\frac{\partial u_{k}({\bf x},t)}{\partial x_{i}}\right)^{2}\;,$ (1)
$\varepsilon_{mag}({\bf x},t)=$ $\frac{\lambda}{2}\sum_{i,k}\left(\frac{\partial h_{i}({\bf x},t)}{\partial x_{k}}+\frac{\partial h_{k}({\bf x},t)}{\partial x_{i}}\right)^{2}\;,$ (2)
where ${\bf u}({\bf x},t)$ denotes the velocity field, ${\bf h}({\bf x},t)=\sqrt{\mu/4\pi\rho}\;{\bf H}({\bf x},t)$ the rescaled magnetic field, $\rho$ the density of the fluid, $\mu$ the permeability, $\nu$ the kinematic viscosity, and $\lambda$ the magnetic diffusivity (see Appendix A for further details). Fig. 1 (a) depicts a two-dimensional cut through the total energy dissipation rate field $\varepsilon_{tot}({\bf x},t)=\varepsilon_{kin}({\bf x},t)+\varepsilon_{mag}({\bf x},t)$ from a direct numerical simulation (DNS) of MHD turbulence (see Tab. 1 for further details). The dissipation field is highly fluctuating and exhibits strong gradients organized in current sheets in contrast to hydrodynamic turbulence Marino et al. (2022); see Fig. 1 (b), where the peaks are more localized, hinting at the presence of vortex tubes Ishihara et al. (2009); Tanaka and Kida (1993).
The solar wind is weakly collisional, which implies that energy cascades from the MHD scales further down to proton and electron scales Matthaeus et al. (2020). Due to the limited resolution of the instruments onboard most spacecraft and the presence of measurement noise, strong gradients of the energy dissipation at kinetic scales can typically not be resolved directly from in situ solar wind observations. Nonetheless, the scale-wise energy transfer at the scales at which the interplanetary plasma has a fluid-like behavior can be estimated indirectly from third-order moments of velocity and magnetic fields Carbone et al. (2009); Marino and Sorriso-Valvo (2023).
Figure 1: Top: Surface plot of a cut through the total local energy dissipation rate, i.e., the sum of Eqs. (1) and (2), obtained from a direct numerical simulation of homogeneous and isotropic MHD turbulence with $1024^{3}$ spatial points (here only an area of $256\times 256$ points is shown) with $\textrm{Re}_{\lambda,kin}$ and $\textrm{Re}_{\lambda,mag}$ and (see main text and Tab. 1 for additional information). The organization of the dissipation field into sheet-like structures is visible and is dominated by a singular sheet peaking at roughly 40 times the mean value. Bottom: Same plot as on top, but obtained from a snapshot of a direct numerical simulation of homogeneous and isotropic hydrodynamic turbulence provided by the Johns Hopkins turbulence database JHTDB (http://turbulence.pha.jhu.edu) with the same spatial resolution $1024^{3}$ and $\textrm{Re}_{\lambda,kin}=418$ (again, only an area of $256\times 256$ points is shown). The dissipation field is organized more randomly than its MHD counterpart.
This can best be illustrated by considering the limit of vanishing magnetic field in the MHD equations, i.e., the Navier-Stokes equation. Under the assumption of homogeneity, isotropy, and statistical stationarity of the flow, the third-order moment of the longitudinal velocity increment $\delta_{r}u=[{\bf u}({\bf x}+{\bf r},t)-{\bf u}({\bf x},t)]\cdot\frac{{\bf r}}{r}$ is related to the averaged kinetic energy dissipation rate according to
$S_{r\,r\,r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)=\left\langle(\delta_{r}u)^{3}\right\rangle=-\frac{4}{5}\left\langle\varepsilon_{kin}\right\rangle r\;,$ (3)
[…]
This article presents new evidence for nonlocal contributions to kinetic and magnetic energy transfer. To this end, we first revisit the derivation of the P&P law Politano and Pouquet (1998a) in Sec. II, obtaining from first principles an additional nonlocal term that accounts for the influence of magnetic fields induced at large scales on fluctuations at small scales as suggested by the Iroshnikov-Kraichnan phenomenology. The importance of this nonlocal term for estimations of turbulent energy transfer based on third-order moments is then assessed by DNS of homogeneous and isotropic MHD turbulence in Sec. III, and the validity of the model proposed is demonstrated in the plasma under study. Although the original P&P law has been verified in numerous DNS of two- and three-dimensional MHD turbulence Mininni and Pouquet (2009); Sorriso-Valvo et al. (2002); Yousef et al. (2007) (see also Marino and Sorriso-Valvo (2023) for further references), here, we provide analytical and numerical evidence that there can be plasma configurations for which considering the nonlocal contributions to the MHD cascade is critical to correctly infer the plasma heating rate through the third-order law approach. These configurations can be representative of heliospheric plasmas, whose highly dynamical state is such that solar wind may at times be characterized by enhanced turbulence and nonlocal couplings of kinetic and magnetic modes, due to magnetic reconnection Cerri et al. (2022), magnetic switchbacks Hernandez et al. (2021); Telloni et al. (2022) or the interaction between fast and slow solar wind streams in the ecliptic Marino et al. (2011).
II Revisiting the P&P law using the invariant theory of MHD turbulence
[…]
$\dot{E}_{tot}(t)=\frac{3}{2}\frac{\textrm{d}}{\textrm{d}t}\left(u_{rms}^{2}+h_{rms}^{2}\right)=-\langle\varepsilon_{tot}\rangle\;,$ (4)
where we imply summation over identical indices and introduce the root mean square velocity and magnetic fields (please see Appendix A for further details on the derivation of Eq. (4)). In his seminal work, Chandrasekhar generalized the work of von Kármán and Howarth de Karman and Howarth (1938) to an invariant theory of homogeneous and isotropic MHD turbulence Chandrasekhar (1951). Here, we consider evolution equations for the velocity and magnetic field correlation tensors, $C_{i\;j}^{\mathbf{u}\mathbf{u}}({\bf r},t)=\left\langle u_{i}u_{j}^{\prime}\right\rangle=\left\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\right\rangle$ and $C_{i\;j}^{\mathbf{h}\mathbf{h}}({\bf r},t)=\left\langle h_{i}h_{j}^{\prime}\right\rangle=\left\langle h_{i}({\bf x},t)h_{j}({\bf x}+{\bf r},t)\right\rangle$, which can be derived from the MHD equations (under the assumption of homogeneity, see Appendix B and C) according to
$\frac{\partial}{\partial t}\left\langle u_{i}u_{j}^{\prime}\right\rangle-$ $2\frac{\partial}{\partial r_{k}}\left\langle(u_{i}u_{k}-h_{i}h_{k})u_{j}^{\prime}\right\rangle=2\nu\nabla_{{\bf r}}^{2}\left\langle u_{i}u_{j}^{\prime}\right\rangle\;,$ (5)
$\frac{\partial}{\partial t}\left\langle h_{i}h_{j}^{\prime}\right\rangle-$ $2\frac{\partial}{\partial r_{k}}\left\langle(h_{i}u_{k}-h_{k}u_{i})h_{j}^{\prime}\right\rangle=2\lambda\nabla_{{\bf r}}^{2}\left\langle h_{i}h_{j}^{\prime}\right\rangle\;.$ (6)
Here, tensors of third order in Eq. (5), i.e., $C_{(ik)j}^{\mathbf{u}\mathbf{u}{\mathbf{u}}}({\bf r},t)=\left\langle u_{i}u_{k}u_{j}^{\prime}\right\rangle$ and $C_{(ik)j}^{\mathbf{h}\mathbf{h}{\mathbf{u}}}({\bf r},t)=\left\langle h_{i}h_{k}u_{j}^{\prime}\right\rangle$ are symmetric in the indices $i$ and $k$, which is represented by the round brackets $(ik)$. By contrast, the third order tensor in Eq. (6) is anti-symmetric in $i$ and $k$, which is denoted by square brackets $[ik]$. It thus admits a different tensorial form than the latter two (see Appendix F for further details), namely,
$A_{[ki]j}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({\bf r},t)=$ $\left\langle(h_{i}u_{k}-h_{k}u_{i})h_{j}^{\prime}\right\rangle$
$=$ $A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}(r,t)\left(\frac{r_{i}}{r}\delta_{jk}-\frac{r_{k}}{r}\delta_{ij}\right)\;.$ (7)
Furthermore, the indices $r$ and $t$ denote the longitudinal and transverse projections, ${\bf u}_{r}=\frac{{\bf r}}{r}\left(\frac{{\bf r}}{r}\cdot{\bf u}\right)$ and ${\bf u}_{t}=-\left(\frac{{\bf r}}{r}\times\left(\frac{{\bf r}}{r}\times{\bf u}\right)\right)$ respectively (the same projections hold for the magnetic field ${\bf h}({\bf x},t)$, see Appendix F for further derivation). Following Chandrasekhar Chandrasekhar (1951) (see Appendix C for further derivations), we obtain evolution equations for the longitudinal velocity and magnetic field correlation functions
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)=$ $\frac{1}{r^{4}}\left[\frac{\partial}{\partial r}r^{4}\left(C_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r,t)-C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)+2\nu\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)\right)\right]\;,$ (8)
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)=$ $-\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)+2\lambda\frac{1}{r^{4}}\frac{\partial}{\partial r}r^{4}\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\;.$ (9)
Eq. (8) is the generalization of the von Kármán-Howarth equation to MHD turbulence and the additional third-order correlation is due to the Lorentz force $C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)$ in the MHD equations. A similar equation also holds for the second-order longitudinal magnetic field correlation function (9) where the third-order correlation is due to the advection by the velocity field and the corresponding transformation of the electric field ${\bf E}^{\prime}={\bf E}+\mu{\bf u}\times{\bf H}$. Similar equations were derived recently in the context of Hall MHD Galtier (2008), where the Hall term modifies such a transformation.
To derive the equivalent of the four-fifths law (18) for MHD turbulence, we introduce the second and third-order longitudinal structure functions
$S_{r\;r}^{{\bf u}{\bf u}}(r,t)=$ $\left\langle\left(\left[{\bf u}({\bf x}+{\bf r},t)-{\bf u}({\bf x},t)\right]\cdot\frac{{\bf r}}{r}\right)^{2}\right\rangle\;,$ (10)
$S_{r\;r}^{{\bf h}{\bf h}}(r,t)=$ $\left\langle\left(\left[{\bf h}({\bf x}+{\bf r},t)-{\bf h}({\bf x},t)\right]\cdot\frac{{\bf r}}{r}\right)^{2}\right\rangle\;,$ (11)
$S_{r\;r\;r}^{{\bf u}{\bf u}{\bf u}}(r,t)=$ $\left\langle\left(\left[{\bf u}({\bf x}+{\bf r},t)-{\bf u}({\bf x},t)\right]\cdot\frac{{\bf r}}{r}\right)^{3}\right\rangle\;.$ (12)
As shown in Appendix H, these structure functions are related to the longitudinal correlation functions as
$S^{{\bf u}{\bf u}}_{r\;r}(r,t)=$ $2[C_{r\;r}^{\mathbf{u}\mathbf{u}}(0,t)-C_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)]\;,$ (13)
$S^{{\bf h}{\bf h}}_{r\;r}(r,t)=$ $2[C_{r\;r}^{\mathbf{h}\mathbf{h}}(0,t)-C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)]\;,$ (14)
$S^{{\bf u}{\bf u}{\bf u}}_{r\;r\;r}(r,t)=$ $6C_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r,t)\;.$ (15)
Furthermore, we evaluate the longitudinal correlation functions for ${r}=0$, which yields
$C_{r\;r}^{\mathbf{u}\mathbf{u}}(0,t)=u_{rms}^{2}\;,\quad\textrm{and}\ \quad C_{r\;r}^{\mathbf{h}\mathbf{h}}(0,t)=h_{rms}^{2}\;.$ (16)
Combining Eqs. (4), (8, (9), and (13-16) thus yields
$\frac{1}{2}\frac{\partial}{\partial t}\left[S^{{\bf u}{\bf u}}_{r\;r}(r,t)+S^{{\bf h}{\bf h}}_{r\;r}(r,t)\right]=$ $-\frac{2}{3}\langle\varepsilon_{tot}\rangle-\frac{1}{r^{4}}\frac{\partial}{\partial r}\left[r^{4}\left(\frac{1}{6}S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r,t)-C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)\right.\right.$
$\left.\left.+\frac{\partial}{\partial r}\left[\nu S_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)+\lambda S_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\right]\right)\right]+\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)\;.$ (17)
In the following, we assume statistical stationarity, which allows us to set the l.h.s of Eq. (17) to zero. Multiplying by $r^{4}$ and integrating from 00 to $r$ yields
$=-\frac{4}{5}\langle\varepsilon_{tot}\rangle r+\frac{\partial}{\partial r}\left[\nu S_{r\;r}^{\mathbf{u}\mathbf{u}}(r)+\lambda S_{r\;r}^{\mathbf{h}\mathbf{h}}(r)\right]\;.$ (18)
In the inertial range, i.e., for $r$ smaller than integral- and larger than dissipation-length scales, we neglect the viscous terms in square brackets and obtain a relation between third-order statistics on the l.h.s. and a term that is proportional to $r$ on the r.h.s. whose magnitude is determined by the total averaged local energy dissipation rate $\langle\varepsilon_{tot}\rangle$ Friedrich et al. (2016). Nonetheless, in contrast to the hydrodynamic case (3) and to the original derivation by Politano and Pouquet Politano and Pouquet (1998a, b) where the energy transfer is purely local in scale $r$, we obtain a scaling law in the MHD inertial range (18) that exhibits a nonlocal term stemming from the magnetic induction equation and is a direct consequence of the anti-symmetric tensorial form (7). As shown in Appendix E, we can re-cast this nonlocal term as
$\textrm{Re}_{\lambda,kin}$ $\textrm{Re}_{\lambda,mag}$ $u_{rms}$ $h_{rms}$ $\langle\varepsilon_{kin}\rangle$ $\langle\varepsilon_{mag}\rangle$ $\nu=\lambda$ $\eta_{kin}$ $\eta_{mag}$ $\lambda_{kin}$ $\lambda_{mag}$ $L_{kin}$ $L_{mag}$ $\textrm{d}x$ N
252 161 2.43 1.40 2.30 5.65 $1.2\cdot 10^{-3}$ $5.2\cdot 10^{-3}$ $4.2\cdot 10^{-3}$ 0.216 0.079 1.350 0.607 $6.1\cdot 10^{-3}$ $1024^{3}$
418 - 0.686 - 0.103 - $1.85\cdot 10^{-4}$ $2.8\cdot 10^{-4}$ 0.113 - 1.364 - $6.1\cdot 10^{-3}$ $1024^{3}$
Table 1: Characteristic parameters of the direct numerical simulations (DNS) of 3D MHD and hydrodynamic turbulence. Taylor-based Reynolds numbers $\textrm{Re}_{\lambda,kin}=u_{rms}\lambda_{kin}/\nu$ and $\textrm{Re}_{\lambda,mag}=h_{rms}\lambda_{mag}/\lambda$, root mean square velocity $u_{rms}=\sqrt{\langle{\bf u}^{2}\rangle/3}$ and magnetic field $h_{rms}=\sqrt{\langle{\bf h}^{2}\rangle/3}$, averaged kinetic and magnetic energy dissipation rates (1 and (2), kinematic viscosity $\nu$ and magnetic diffusivity $\lambda$, Kolmogorov dissipation length scales $\eta_{kin}=(\nu^{3}/\langle\varepsilon_{kin}\rangle)^{1/4}$ and $\eta_{mag}=(\lambda^{3}/\langle\varepsilon_{mag}\rangle)^{1/4}$, Taylor length scales $\lambda_{kin}=\sqrt{15\nu u_{rms}^{2}/\langle\varepsilon_{kin}\rangle}$ and $\lambda_{mag}=\sqrt{15\lambda h_{rms}^{2}/\langle\varepsilon_{mag}\rangle}$, and integral length scales $L_{kin}$ and $L_{mag}$ (both determined from the correlation functions), grid spacing $\textrm{d}x$, and number of grid points $N$. Statistical quantities in the MHD and the hydrodynamic turbulence simulation were averaged over approximately 9 and 5 large eddy turnover times, respectively.
$\int_{0}^{r}\textrm{d}r^{\prime}r^{\prime 3}A_{[rt]\;t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r^{\prime})=-\int_{r}^{\infty}\textrm{d}r^{\prime}r^{\prime 3}A_{[rt]\;t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r^{\prime})\;,$ (19)
which thus represents the influence of magnetic field structures induced at larger scales $r^{\prime}$ on the local fluctuations at scale $r$ in the sense of the Alfvén effect Alfvén (1950); Biskamp (2003); Iroshnikov (1964); Kraichnan (1965). We will now consider further implications of this additional nonlocality for the conservation of energy at small scales and the phenomenon of solar wind heating by turbulent cascades. In more detail, for small scales $r$ in the vicinity of the dissipation range, nonlinear transfer terms on the l.h.s. in Eq. (18) should decay faster than the terms on the r.h.s. Monin and Yaglom (2007). From a Taylor expansion around $r=0$, we obtain $S_{r\;r\,r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)\sim r^{3}$ as well as
[…]
Figure 2: Temporal evolution of the kinetic, magnetic, and total energy dissipation rates throughout the simulation. After an initial oscillation, the dissipation rates remain fairly constant due to the applied forcing mechanism, which indicates a close-to-stationary MHD flow. The assessed statistical quantities in Tab. 1 were averaged for $t/T_{L}>5$.
Figure 3: (a) Assessment of the four-fifths law in MHD turbulence from DNS ($1024^{3}$) of forced MHD turbulence simulation. The dashed line represents the r.h.s. of Eq. (18, which is determined from the total energy dissipation rate $\varepsilon_{tot}$ of the snapshot. The top curve corresponds the original P&P law $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)-6C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r)$ derived in Politano and Pouquet (1998a, a) whereas the additional nonlocal term in Eq. (18) is closer to the dashed line. The mixed tensors verify the dissipation range behavior (22). The lowest curve represents the third-order longitudinal velocity structure function $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)$. (b) Same as in (a) but compensated by the r.h.s. in Eq. (18). The original P&P law (top curve) overestimates the total energy dissipation rate $\varepsilon_{tot}$ by a factor of 1.71 and peaks in front of the inertial range, whereas the inclusion of the nonlocal term in Eq. (18) leads to the correct prediction (dashed line).
Hence, for small $r$, the third-order velocity contribution can be neglected. In contrast, the two other correlation functions (20) and (21) would yield a finite contribution to the nonlinear transfer in the dissipation range (please note that the l.h.s in Eq. (18) scales as $r$ as well). As was first derived by Chandrasekhar, for small $r$, both correlation tensors are related by
$\left.\frac{\partial A_{[rt]r}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r)}{\partial r}\right|_{r=0}=-\frac{5}{4}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r)}{\partial r}\right|_{r=0}\;,$ (22)
which is derived in Appendix D. Inserting this relation into Eq. (18) leads to the exact cancelation of both third-order correlations and reproduces the expected result in the dissipation range, i.e.,
$0=-\frac{4}{5}\langle\varepsilon_{tot}\rangle r+\frac{\partial}{\partial r}\left[\nu S_{r\;r}^{\mathbf{u}\mathbf{u}}(r)+\lambda S_{r\;r}^{\mathbf{h}\mathbf{h}}(r)\right]\;,$ (23)
or
$S_{r\;r}^{\mathbf{h}\mathbf{h}}(r)+\textrm{Pm}S_{r\;r}^{\mathbf{u}\mathbf{u}}(r)=\frac{\langle\varepsilon_{tot}\rangle}{15\lambda}r^{2}\;,$ (24)
where $\textrm{Pm}={\nu}/{\lambda}$ denotes the magnetic Prandtl number. In the following section, we will verify the modified scaling law (18) in direct numerical simulations of MHD turbulence.
III Verification of the third-order law via direct numerical simulations of MHD turbulence
To verify Eq. (18), we performed DNS of homogeneous and isotropic MHD turbulence in a box of $1024^{3}$ grid points with characteristic turbulence parameters summarized in Tab. 1. The forcing scheme consists of a divergence-free random forcing Alvelius (1999) that is applied to the evolution equation of the velocity field (i.e., the evolution equation for the magnetic field is not actively forced). As shown in Fig. 2, after an initial phase, the forcing scheme leads to nearly constant kinetic and magnetic energy dissipation rates $\langle\varepsilon_{kin}\rangle$ and $\langle\varepsilon_{mag}\rangle$ over time. Fig. 3 depicts the third-order moments that enter Eq. (18). The lowest curve corresponds to the third-order velocity structure function $S_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r)$. Interestingly, this curve exhibits three distinct scaling regions: In the dissipation range, we observe the predicted scaling $\sim r^{3}$ followed by a region that scales in agreement with the Iroshnikov-Kraichnan phenomenology $\sim r^{3/4}$ whereas a linear increase seems to dominate at larger $r$. Furthermore, in the dissipation range, we can directly verify the relation between the symmetric and anti-symmetric correlations (22). In the case under study, the original prediction Politano and Pouquet (1998a, b) without the additional nonlocal source term overestimates the total energy dissipation rate $\langle\varepsilon_{tot}\rangle$ (dashed line) and does not vanish for small $r$, therefore, violating energy conservation. This becomes even more apparent in the compensated plot in Fig. 3 (b), which suggests that including the source term results in the correct scaling in the inertial range in between the largest Taylor scale $\lambda_{kin}$ and the smallest integral length scale $L_{mag}$. In contrast, the scaling obtained using the P&P law Politano and Pouquet (1998a, b) (green curve) would plateau before this inertial range and overestimate the $\langle\varepsilon_{tot}\rangle$ by a factor of 1.71. We must emphasize that this discrepancy between the original P&P law and (18) might have been barely detectable in previous numerical studies of MHD turbulence Mininni and Pouquet (2009); Sorriso-Valvo et al. (2002); Yousef et al. (2007) due for instance to low magnetic and fluid Reynolds numbers and signatures of the large-scale anisotropic forcing scheme (e.g., by Taylor-Green vortices). Indeed, MHD and kinetic simulations reproduce plasma dynamics and features peculiar to the set of governing parameters defined for specific runs, which may differ significantly from actual solar wind parameters. Due to the insufficient computational power of today’s supercomputers, constraints in reproducing the solar wind numerically are technical but also inherent since solar wind dynamics develop in a vast parameter space, hardly reproducible even with large ensembles of simulations.
This entails that the influence of the additional nonlocal term on cascade processes can be sub-dominant to the third-order correlation stemming from the Lorentz force $C_{r\;r\;r}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)$ or that the assumptions of isotropy and homogeneity in the derivation of Eq. (18) are not strictly guaranteed, in some of the simulations appearing in the literature, and also in many cases within the actual solar wind. Nonetheless, for some settings, e.g., strong Alfvénic turbulence, this term might play an important, non-negligible role.
IV Conclusions
In this paper, we have derived from first principles an additional nonlocal term in the equation for the energy cascade directly from the incompressible MHD equations, assuming isotropy and homogeneity only. This additional term can potentially improve the heating rate predictions by the original P&P law when nonlocal interactions become essential in the overall plasma dynamics. We have verified by numerical simulations of MHD turbulence forced at large scales that this additional nonlocal term is critical in describing the energy transfer from large to small scales for the set of parameters chosen.
The exact law established by Politano and Pouquet Politano and Pouquet (1998a, b) assumes homogeneity, isotropy, and incompressibility - conditions that are fulfilled perhaps only locally in interplanetary space plasmas - representing to date the most robust statistical framework to characterize energy transfer and heating in the solar wind Marino and Sorriso-Valvo (2023); Marino et al. (2008). Without diminishing its validity, our findings suggest that in some solar wind samples, the current analyses of observational data by the original P&P law and some of its surrogates might substantially overestimate the rate at which energy is dissipated (or further transferred across sub-proton scales) at the bottom of the MHD turbulent cascade developing in the solar wind. This leads to improved third-order moment law predictions characterizing the observed space plasma heating.
As mentioned in Sec. II, the additional nonlocal term arises due to the advection of the magnetic field by the velocity field and the corresponding transformation of the electric field ${\bf E}^{\prime}={\bf E}+\mu{\bf u}\times{\bf H}$. Hence, it would also be interesting to assess other transformations of the electric field, e.g., in the context of Hall MHD Galtier (2008); Foldes et al. (2023). Future work will be devoted to applying the law (18) and assessing the influence of nonlocality in solar wind measurements. Moreover, we aim at reconstructing velocity and magnetic fields from partial measurements by extending current velocity field reconstructions Friedrich et al. (2022); Lübke et al. (2023) to MHD turbulence Friedrich (2020); Friedrich and Grauer (2020); Lübke et al. (2024), exploiting observations of state-of-art ongoing (Owen et al., 2020; Fox et al., 2016) and future multi-spacecraft solar wind mission (Klein et al., 2023; Pecora et al., 2023a, b; Retino et al., 2022).
V Acknowledgments
J.F. acknowledges fruitful discussions with H. Homann, A. Pumir, and J. Peinke. We gratefully acknowledge the Gauss Centre for Supercomputing (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS Alvarez (2021) at Jülich Supercomputing Centre (JSC). The simulation output was also analyzed on HPC facilities at the École Centrale de Lyon (PMCS2I) in Ecully (France) and the HPC Cluster STORM, located at the University of Oldenburg (Germany) and funded through the REACT-EU program by the NBank and the Ministry of Science and Culture (MWK) of the Lower Saxony State under grant number ZW7-85186744. R.M. acknowledges support from the project “EVENTFUL” (ANR-20-CE30-0011), funded by the French “Agence Nationale de la Recherche” - ANR through the program AAPG-2020. M.W.  acknowledges support from the German Science Foundation DFG within the Collaborative Research Center SFB1491.
Appendix A Derivation of the Equation of Total Energy Balance in MHD Turbulence
In this section, we derive an evolution equation for the total energy (i.e., the sum of kinetic and magnetic energy) in MHD turbulence. Therefore, we consider the MHD equations
$\frac{\partial}{\partial t}u_{i}({\bf x},t)+u_{k}({\bf x},t)\frac{\partial}{\partial x_{k}}u_{i}({\bf x},t)-h_{k}({\bf x},t)\frac{\partial}{\partial x_{k}}h_{i}({\bf x},t)$ $=-\frac{1}{\rho}\frac{\partial}{\partial x_{i}}P({\bf x},t)+\nu\nabla^{2}_{\bf x}u_{i}({\bf x},t)\;,$ (25)
$\frac{\partial}{\partial t}h_{i}({\bf x},t)+u_{k}({\bf x},t)\frac{\partial}{\partial x_{k}}h_{i}({\bf x},t)-h_{k}({\bf x},t)\frac{\partial}{\partial x_{k}}u_{i}({\bf x},t)$ $=\lambda\nabla^{2}_{\bf x}h_{i}({\bf x},t)\;,$ (26)
where summation over identical indices is implied. Here, $P({\bf x},t)=p({\bf x},t)+\frac{\rho}{2}|{\bf h}|^{2}$ denotes the sum of hydrodynamic pressure $p({\bf x},t)$ and the magnetic pressure, $\rho$ the density (which will be set to one), $\nu$ the kinematic viscosity and $\lambda$ the magnetic diffusivity of the conducting fluid. Furthermore, it should be noted that in this formulation of the MHD equations, the magnetic field ${\bf H}$ is measured in units of the velocity field
${\bf h}({\bf x},t)=\sqrt{\frac{\mu}{4\pi\rho}}{\bf H}({\bf x},t)\;.$ (27)
The total energy can now be defined as the sum of kinetic and magnetic energy according to
$E_{tot}(t)=\frac{1}{2}\left\langle u_{i}u_{i}+h_{i}h_{i}\right\rangle=\frac{3}{2}\left(u_{rms}^{2}+h_{rms}^{2}\right)\;,$ (28)
[…]
$\ \frac{1}{2}\frac{\partial}{\partial t}{\bf u}^{2}+\frac{1}{2}\frac{\partial}{\partial x_{k}}u_{k}{\bf u}^{2}-u_{i}\frac{\partial}{\partial x_{k}}h_{k}h_{i}=$ $-\frac{1}{\rho}\frac{\partial}{\partial x_{i}}u_{i}P$ (29)
$+u_{i}\nu\nabla^{2}_{\bf x}u_{i}\;,$
$\frac{1}{2}\frac{\partial}{\partial t}{\bf h}^{2}+\frac{1}{2}\frac{\partial}{\partial x_{k}}u_{k}{\bf h}^{2}-h_{i}\frac{\partial}{\partial x_{k}}h_{k}u_{i}=$ $h_{i}\lambda\nabla^{2}_{\bf x}h_{i}\;,$ (30)
where we used the incompressibility conditions for both velocity and magnetic field, i.e., $\frac{\partial}{\partial x_{k}}u_{k}=0$ and $\frac{\partial}{\partial x_{k}}h_{k}=0$. Here, the viscous terms can be reformulated according to
$=$ $\frac{\nu}{2}\nabla^{2}_{\bf x}{\bf u}^{2}-\frac{\nu}{2}\sum_{i,k}\left(\frac{\partial u_{i}}{\partial x_{k}}+\frac{\partial u_{k}}{\partial x_{i}}\right)^{2}+\nu\sum_{i,k}\left(\frac{\partial u_{i}}{\partial x_{k}}\frac{\partial u_{k}}{\partial x_{i}}\right)$
$=$ $\frac{\nu}{2}\nabla^{2}_{\bf x}{\bf u}^{2}-\varepsilon_{kin}+\nu\frac{\partial}{\partial x_{i}}u_{k}\frac{\partial}{\partial x_{k}}u_{i}\;.$ (31)
The same procedure applies to the diffusive terms in Eq. (30). Furthermore, local kinetic and magnetic energy dissipation rates are defined according to
$\varepsilon_{kin}({\bf x},t)=$ $\frac{\nu}{2}\sum_{i,k}\left(\frac{\partial u_{i}({\bf x},t)}{\partial x_{k}}+\frac{\partial u_{k}({\bf x},t)}{\partial x_{i}}\right)^{2}\;,$ (32)
$\varepsilon_{mag}({\bf x},t)=$ $\frac{\lambda}{2}\sum_{i,k}\left(\frac{\partial h_{i}({\bf x},t)}{\partial x_{k}}+\frac{\partial h_{k}({\bf x},t)}{\partial x_{i}}\right)^{2}\;.$ (33)
We observe that nonlinear terms which are advected by the magnetic field (third terms on l.h.s of Eqs. (29) and (30)) can be cast in conservative form by adding Eqs. (29) and (30), which yields
$\frac{\partial}{\partial t}e_{tot}({\bf x},t)+\nabla_{{\bf x}}\cdot{\bf J}^{tot}({\bf x},t)=-\varepsilon_{tot}({\bf x},t)\;,$ (34)
where we defined
$e_{tot}({\bf x},t)=\frac{{\bf u}^{2}({\bf x},t)+{\bf h}^{2}({\bf x},t)}{2}\;,$ (35)
${\bf J}^{tot}({\bf x},t)={\bf u}({\bf x},t)\left(\frac{{\bf u}^{2}({\bf x},t)+{\bf h}^{2}({\bf x},t)}{2}+\frac{P({\bf x},t)}{\rho}\right)$
$-{\bf h}({\bf x},t)[{\bf u}({\bf x},t)\cdot{\bf h}({\bf x},t)]-\frac{\nu}{2}\nabla_{\bf x}{\bf u}^{2}({\bf x},t)-\frac{\lambda}{2}\nabla_{\bf x}{\bf h}^{2}({\bf x},t)$
$-\nu{\bf u}({\bf x},t)\cdot\nabla_{{\bf x}}{\bf u}({\bf x},t)-\lambda{\bf h}({\bf x},t)\cdot\nabla_{{\bf x}}{\bf h}({\bf x},t)\;,$ (36)
$\varepsilon_{tot}({\bf x},t)=\varepsilon_{kin}({\bf x},t)+\varepsilon_{mag}({\bf x},t)\;.$ (37)
The two first terms in the brackets of (A denote the kinetic and magnetic energy transported throughout the surface. The second term is the work from the fluid against the total pressure. The next term is the cross helicity ${\bf u}({\bf x},t)\cdot{\bf h}({\bf x},t)$ which is transported anti-parallel along the lines of force. Performing the averaging procedure, we obtain a temporal evolution equation for the total energy $E_{tot}(t)=\langle e_{tot}({\bf x},t)\rangle$ as
$\dot{E}_{tot}(t)=-\langle\varepsilon_{tot}({\bf x},t)\rangle\;,$ (38)
where we assumed homogeneity to neglect the transport terms $\nabla_{\bf x}\cdot\langle{\bf J}^{tot}({\bf x},t)\rangle=0$. Hence, the total energy is only changed by viscous and magnetic dissipation. Here, we only considered decaying MHD turbulence; otherwise, the temporal evolution $E_{tot}(t)$ would be influenced by additional forcing mechanisms in Eqs. (25) and (26) that would lead to additional source terms on the r.h.s. of Eq. (38).
Appendix B Derivation of the Friedmann-Keller Hierarchy in Homogeneous and Isotropic MHD Turbulence
In this section, we recapitulate Chandrasekhar’s derivation Chandrasekhar (1951) of the Friedmann-Keller hierarchy Keller and Friedman (1924) for magnetic and velocity field correlation functions. In a statistical description of MHD turbulence, evolution equations for the two-point correlation tensors
$C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)=$ $\left\langle u_{i}({\bf x},t)u_{j}({\bf x}^{\prime},t)\right\rangle\;,$ (39)
$C_{i\;j}^{{\bf u}{\bf h}}({\bf r},t)=$ $\left\langle u_{i}({\bf x},t)h_{j}({\bf x}^{\prime},t)\right\rangle\;,$ (40)
$C_{i\;j}^{{\bf h}{\bf h}}({\bf r},t)=$ $\left\langle h_{i}({\bf x},t)h_{j}({\bf x}^{\prime},t)\right\rangle\;,$ (41)
are obtained by the same procedure as in the hydrodynamic case de Karman and Howarth (1938) and were first derived by Chandrasekhar Chandrasekhar (1951). In the following, we are solely interested in evolution equations for the two-point velocity and magnetic field tensors (39) and (41). The evolution equation and invariant form of the cross helicity tensor (40), which is a skew-symmetric tensor, can be found in Chandrasekhar (1951). The procedure for the velocity correlation tensor (39) starts with multiplying Eq. (25) by $u_{j}^{\prime}=u_{j}({\bf x}^{\prime},t)$
$=-u_{j}^{\prime}\frac{1}{\rho}\frac{\partial}{\partial x_{i}}P+\nu u_{j}^{\prime}\nabla^{2}_{\bf x}u_{i}\;,$ (42)
where we used the incompressibility condition for both velocity and magnetic field, i.e., $\frac{\partial}{\partial x_{k}}u_{k}=0$ and $\frac{\partial}{\partial x_{k}}h_{k}=0$. In the same manner, we can multiply the evolution equation for $u_{j}({\bf x}^{\prime},t)$ by $u_{i}=u_{i}({\bf x},t)$, which yields
$=-u_{i}\frac{1}{\rho}\frac{\partial}{\partial x_{i}^{\prime}}P^{\prime}+\nu u_{i}\nabla^{2}_{{\bf x}^{\prime}}u_{j}^{\prime}\;.$ (43)
Moreover, from the induction equation (26), we obtain
$h_{j}^{\prime}\frac{\partial}{\partial t}h_{i}+\frac{\partial}{\partial x_{k}}\left(u_{k}h_{i}h_{j}^{\prime}-h_{k}u_{i}h_{j}^{\prime}\right)=\lambda h_{j}^{\prime}\nabla^{2}_{\bf x}h_{i}\;,$ (44)
as well as
$h_{i}\frac{\partial}{\partial t}h_{j}^{\prime}+\frac{\partial}{\partial x_{k}^{\prime}}\left(u_{k}^{\prime}h_{j}^{\prime}h_{i}-h_{k}^{\prime}u_{j}^{\prime}h_{i}\right)=\lambda h_{i}\nabla^{2}_{{\bf x}^{\prime}}h_{j}^{\prime}\;.$ (45)
[…]
  • Based on the assumption of homogeneity, correlation functions solely depend on the relative distance ${\bf r}={\bf x}^{\prime}-{\bf x}$, and we obtain
    $\quad C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)=\langle u_{i}({\bf x},t)u_{j}({\bf x}^{\prime},t)\rangle=C_{i\;j}^{{\bf u}{\bf u}}(-{\bf r},t)\;,$ (46)
    $\quad C_{(ki)j}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)=\langle u_{k}({\bf x},t)u_{i}({\bf x},t)u_{j}({\bf x}^{\prime},t)\rangle\;,$ (47)
    $\quad C_{(kj)i}^{{\bf u}{\bf u}{\bf u}}(-{\bf r},t)=\langle u_{k}({\bf x}^{\prime},t)u_{j}({\bf x}^{\prime},t)u_{i}({\bf x},t)\rangle\;.$ (48)
    Therefore, viscous terms can be rewritten according to
    $=$ $2\nabla^{2}_{\bf r}\langle u_{i}({\bf x},t)u_{j}({\bf x}^{\prime},t)\rangle=2\nabla^{2}_{\bf r}C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)\;.$ (49)
    Correlations where the magnetic field occurs an even number of times transform identically to Eqs. (46-48)
    $\;\;\;\quad C_{i\;j}^{{\bf h}{\bf h}}({\bf r},t)=\langle h_{i}({\bf x},t)h_{j}({\bf x}^{\prime},t)\rangle=C_{i\;j}^{{\bf h}{\bf h}}(-{\bf r},t)\;,$ (50)
    $\;\;\;\quad C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)=\langle h_{k}({\bf x},t)h_{i}({\bf x},t)u_{j}({\bf x}^{\prime},t)\rangle\;,$ (51)
    $\;\;\;\quad C_{(kj)i}^{{\bf h}{\bf h}{\bf u}}(-{\bf r},t)=\langle h_{k}({\bf x}^{\prime},t)h_{j}({\bf x}^{\prime},t)u_{i}({\bf x},t)\rangle\;,$ (52)
    and for the anti-symmetric tensor, we obtain
    $A_{[ki]j}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)$ (53)
    $\quad$== $\langle(u_{k}({\bf x},t)h_{i}({\bf x},t)-u_{i}({\bf x},t)h_{k}({\bf x},t))h_{j}({\bf x}^{\prime},t)\rangle\;,$
    $A_{[kj]i}^{{\bf u}{\bf h}{\bf h}}(-{\bf r},t)$ (54)
    $\quad$== $\langle(u_{k}({\bf x}^{\prime},t)h_{j}({\bf x}^{\prime},t)-u_{j}({\bf x}^{\prime},t)h_{k}({\bf x}^{\prime},t))h_{i}({\bf x},t)\rangle\;.$
    Here, square brackets in the index of Eq. (53) indicate that the corresponding tensor is anti-symmetric in $i$ and $k$.
  • Furthermore, isotropic and mirror-symmetric tensors of the third order obey the following relation
    $C_{(kj)i}^{{\bf u}{\bf u}{\bf u}}(-{\bf r},t)=$ $-C_{(kj)i}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)\;,$ (55)
    $C_{(kj)i}^{{\bf h}{\bf h}{\bf u}}(-{\bf r},t)=$ $-C_{(kj)i}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)\;,$ (56)
    and for the anti-symmetric tensor
    $A_{[kj]i}^{{\bf u}{\bf h}{\bf h}}(-{\bf r},t)=-A_{[kj]i}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)\;.$ (57)
  • Pressure-velocity contributions vanish based on isotropy and mirror symmetry Faust et al. (2015), which yields
    $\left(\frac{\partial}{\partial x_{i}}+\frac{\partial}{\partial x_{i}^{\prime}}\right)\langle P({\bf x},t)u_{j}({\bf x}^{\prime},t)\rangle=0\;.$ (58)
Under the above assumptions, we obtain from Eqs. (42) and (43)
$\frac{\partial}{\partial t}C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)-\frac{\partial}{\partial r_{k}}\left(C_{(ki)j}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)+C_{(kj)i}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)\right.$
$\left.-C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)-C_{(kj)i}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)\right)=2\nu\nabla^{2}_{\bf r}C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)\;.$ (59)
Furthermore, as terms like $\frac{\partial}{\partial r_{k}}C_{(ki)j}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)$ are second order isotropic tensors, they must be symmetric in $i$ and $j$. We thus obtain
$\frac{\partial}{\partial t}C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)-2\frac{\partial}{\partial r_{k}}\left(C_{(ki)j}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)-C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)\right)$
$=$ $2\nu\nabla^{2}_{\bf r}C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)\;.$ (60)
The same procedure can be applied to the induction equation (26) and we obtain
$\frac{\partial}{\partial t}C_{i\;j}^{{\bf h}{\bf h}}({\bf r},t)-2\frac{\partial}{\partial r_{k}}A_{[ki]j}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)=2\lambda\nabla^{2}_{\bf r}C_{i\;j}^{{\bf h}{\bf h}}({\bf r},t)\;.$ (61)
These two equations (and the additional equation for the cross helicity tensor of second order derived by Chandrasekhar Chandrasekhar (1951)) are a generalization of the Friedmann-Keller hierarchy Keller and Friedman (1924) to homogeneous and isotropic MHD turbulence.
Appendix C Derivation of Evolution Equations for Longitudinal Velocity and Magnetic Field Correlation Functions of Second Order
A further simplification of Eqs. (60) and (61) can be obtained from the invariant theory of homogeneous and isotropic turbulence Robertson (1940). Each tensor can be rewritten in terms of its longitudinal correlation function, which is a consequence of the incompressibility condition and is further derived in Appendices F,G, and H. The second-order tensors (46) and (50) both follow the tensorial form
$C_{ij}({\bf r},t)=$ $\left(C_{rr}(r,t)-\frac{1}{2r}\frac{\partial}{\partial r}(r^{2}C_{rr}(r,t))\right)\frac{r_{i}r_{j}}{r^{2}}$
$+\frac{1}{2r}\frac{\partial}{\partial r}(r^{2}C_{rr}(r,t))\delta_{ij}\;,$ (62)
where $C_{rr}(r,t)$ denotes the longitudinal correlation function of second order. In the same manner, third-order tensors that are symmetric in $i$ and $k$, such as (48) and (51) obey the following form
$C_{(ki)j}({\bf r},t)=-\frac{r^{2}}{2}\frac{\partial}{\partial r}\left(\frac{C_{rrr}(r,t)}{r}\right)\frac{r_{i}r_{j}r_{k}}{r^{3}}$
$+\frac{1}{4r}\frac{\partial}{\partial r}\left(r^{2}C_{rrr}(r,t)\right)\left(\frac{r_{i}}{r}\delta_{kj}+\frac{r_{k}}{r}\delta_{ij}\right)$
$-\frac{C_{rrr}(r,t)}{2}\frac{r_{j}}{r}\delta_{ik}\;.$ (63)
It should be noted that this tensorial form is only applicable to third-order terms in Eq. (60). The third-order correlation in Eq. (61) is anti-symmetric in $k$ and $i$ and follows a different tensorial form namely
$A_{[ki]j}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({\bf r},t)=A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}(r,t)\left(\frac{r_{i}}{r}\delta_{jk}-\frac{r_{k}}{r}\delta_{ij}\right)\;.$ (64)
Summing over $i=j$ in Eqs. (60-61) yields
$Q_{kin}(r,t)$ $=$ $\frac{1}{2}\sum_{i=j}C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)\;,\quad J^{kin}_{k}({\bf r},t)=-\sum_{i=j}\left(C_{(ki)j}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)-C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)\right)\;,$ (65)
$Q_{mag}(r,t)$ $=$ $\frac{1}{2}\sum_{i=j}C_{ij}({\bf r},t)\;,\quad J^{mag}_{k}({\bf r},t)=-\sum_{i=j}A_{[ki]j}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)\;,$ (66)
and we obtain two balance equation for $Q_{kin}(r,t)$ and $Q_{mag}(r,t)$ with their corresponding currents ${\bf J}^{kin}({\bf r},t)$ and ${\bf J}^{mag}({\bf r},t)$ that read
$\frac{\partial}{\partial t}Q_{kin}(r,t)+\nabla_{\bf r}\cdot{\bf J}^{kin}({\bf r},t)=$ $2{\nu}\Delta_{\bf r}Q_{kin}(r,t)\;,$ (67)
$\frac{\partial}{\partial t}Q_{mag}(r,t)+\nabla_{\bf r}\cdot{\bf J}^{mag}({\bf r},t)=$ $2{\lambda}\Delta_{\bf r}Q_{mag}(r,t)\;.$ (68)
As shown in Appendix H, $Q_{kin}(r,t)$, $Q_{mag}(r,t)$ and their corresponding currents ${\bf J}^{kin}({\bf r},t)$, ${\bf J}^{mag}({\bf r},t)$ can be expressed as
$\frac{\partial}{\partial t}Q_{kin}(r,t)$ $=$ $\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(r^{3}\frac{\partial}{\partial t}C_{r\;r}^{{\bf u}{\bf u}}(r,t)\right),\qquad\frac{\partial}{\partial t}Q_{mag}(r,t)=\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(r^{3}\frac{\partial}{\partial t}C_{r\;r}^{{\bf h}{\bf h}}(r,t)\right),$ (69)
$\frac{\partial}{\partial r_{k}}J_{k}^{kin}({\bf r},t)$ $=$ $-\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(\frac{1}{r}\frac{\partial}{\partial r}\left[r^{4}\left(C_{r\;r\;r}^{{\bf u}{\bf u}{\bf u}}-C_{r\;r\;r}^{{\bf h}{\bf h}{\bf u}}\right)\right]\right)\;,\qquad\frac{\partial}{\partial r_{k}}J_{k}^{mag}({\bf r},t)=\frac{2}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}(r,t)\right)\;,$ (70)
$\Delta_{\bf r}Q_{kin}(r,t)$ $=$ $\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}\frac{\partial}{\partial r}Q_{kin}(r,t)\right)\;,\qquad\Delta_{\bf r}Q_{mag}(r,t)=\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}\frac{\partial}{\partial r}Q_{mag}(r,t)\right)\;.$ (71)
Inserting these relations into Eq. (67) and (68) yields
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)=$ $\frac{1}{r^{4}}\Bigg{[}\frac{\partial}{\partial r}r^{4}\Bigg{(}C_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r,t)-C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)$
$+2\nu\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)\Bigg{)}\Bigg{]}\;,$ (72)
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)=$ $-\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)+2\lambda\frac{1}{r^{4}}\frac{\partial}{\partial r}r^{4}\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\;.$ (73)
Eq. (72) is the generalization of the von Kármán-Howarth equation to MHD turbulence with the additional third-order correlation function $C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)$ that is due to the Lorentz force in Eq. (25). The evolution equation for the second-order longitudinal correlation function for the magnetic field (73) possesses a different differential form, which is due to the anti-symmetry of the tensor in Eq. (64).
Appendix D On the Equivalence of the Mixed Third-Order Correlations for Small Scale Separations
As first recognized by Chandrasekhar Chandrasekhar (1951), the defining scalars of the mixed third-order correlation functions $C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)=\left\langle h_{i}({\bf x},t)h_{k}({\bf x},t)u_{j}({\bf x}^{\prime},t)\right\rangle$ and $A_{[ki]j}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)=\left\langle(h_{i}({\bf x},t)u_{k}({\bf x},t)-h_{k}({\bf x},t)u_{i}({\bf x},t))h_{j}({\bf x}^{\prime},t)\right\rangle$ are related to each other for small scale separations ${\bf r}$. This can be shown as follows: First, we consider the derivative of the symmetric tensor defined by Eq. (63)
$\left.\frac{\partial C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)}{\partial r_{i}}\right|_{{\bf r}=0}=\left\langle h_{i}({\bf x},t)h_{k}({\bf x},t)\frac{\partial u_{j}({\bf x},t)}{\partial x_{i}}\right\rangle$
$=$ $\left[\frac{C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{r}+\frac{3}{2}\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right]_{r=0}\delta_{jk}$
$=$ $\frac{5}{2}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right|_{r=0}\delta_{jk}\;.$ (74)
Similarly, we derive the tensorial form of the anti-symmetric third-order correlation tensor (64) and obtain
$\left.\frac{\partial A_{[ki]j}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)}{\partial r_{i}}\right|_{{\bf r}=0}$
$=\left\langle(h_{i}({\bf x},t)u_{k}({\bf x},t)-h_{k}({\bf x},t)u_{i}({\bf x},t))\frac{\partial h_{j}({\bf x},t)}{\partial x_{i}}\right\rangle$
$=$ $\left.2\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}\delta_{jk}\;.$ (75)
Next, we derive the following identity
$=$ $\left\langle h_{i}({\bf x},t)u_{k}({\bf x},t)\frac{\partial h_{j}({\bf x},t)}{\partial x_{i}}\right\rangle-\left\langle h_{k}({\bf x},t)u_{i}({\bf x},t)\frac{\partial h_{j}({\bf x},t)}{\partial x_{i}}\right\rangle$
$=$ $\underbrace{\frac{\partial}{\partial x_{i}}\left\langle h_{i}({\bf x},t)h_{j}({\bf x},t)u_{k}({\bf x},t)\right\rangle}_{=0,\;\textrm{homogeneity}}-\left\langle h_{i}({\bf x},t)h_{j}({\bf x},t)\frac{\partial u_{k}({\bf x},t)}{\partial x_{i}}\right\rangle-\left\langle\underbrace{\frac{\partial h_{i}({\bf x},t)}{\partial x_{i}}}_{=0}h_{j}({\bf x},t)u_{k}({\bf x},t)\right\rangle$
$-\underbrace{\frac{\partial}{\partial x_{i}}\left\langle h_{k}({\bf x},t)u_{i}({\bf x},t)h_{j}({\bf x},t)\right\rangle}_{=0,\;\textrm{homogeneity}}+\left\langle h_{j}({\bf x},t)u_{i}({\bf x},t)\frac{\partial h_{k}({\bf x},t)}{\partial x_{i}}\right\rangle+\left\langle\underbrace{\frac{\partial u_{i}({\bf x},t)}{\partial x_{i}}}_{=0}h_{j}({\bf x},t)h_{k}({\bf x},t)\right\rangle$
$=$ $-\left\langle h_{i}({\bf x},t)h_{j}({\bf x},t)\frac{\partial u_{k}({\bf x},t)}{\partial x_{i}}\right\rangle+\left\langle h_{j}({\bf x},t)u_{i}({\bf x},t)\frac{\partial h_{k}({\bf x},t)}{\partial x_{i}}\right\rangle\;.$ (76)
Interchanging $j$ and $k$ in Eq. (75) yields
$\left\langle(h_{i}({\bf x},t)u_{j}({\bf x},t)-h_{j}({\bf x},t)u_{i}({\bf x},t))\frac{\partial h_{k}({\bf x},t)}{\partial x_{i}}\right\rangle$
$=$ $\underbrace{\frac{\partial}{\partial x_{i}}\left\langle h_{i}({\bf x},t)h_{k}({\bf x},t)u_{j}({\bf x},t)\right\rangle}_{=0,\;\textrm{homogeneity}}$
$-\left\langle h_{i}({\bf x},t)h_{k}({\bf x},t)\frac{\partial u_{j}({\bf x},t)}{\partial x_{i}}\right\rangle$
$-\left\langle\underbrace{\frac{\partial h_{i}({\bf x},t)}{\partial x_{i}}}_{=0}h_{k}({\bf x},t)u_{j}({\bf x},t)\right\rangle$
$-\left\langle h_{j}({\bf x},t)u_{i}({\bf x},t)\frac{\partial h_{k}({\bf x},t)}{\partial x_{i}}\right\rangle\;.$ (77)
Adding Eqs. (76) and (77) yields
$4\left.\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}\delta_{jk}=-\left\langle h_{i}({\bf x},t)h_{j}({\bf x},t)\frac{\partial u_{k}({\bf x},t)}{\partial x_{i}}\right\rangle$
[…]
Hence, for small-scale separations $r$, the defining scalars of symmetric and anti-symmetric mixed third-order tensors are related according to
$\left.\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}=-\frac{5}{4}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right|_{r=0}\;.$ (79)
Appendix E Loitsiansky Invariants and their Implications for the Additional Source Term
In this section, we discuss further implications of Eqs. (8) and (9), namely the existence of certain invariants. As shown by Loitsiansky Loitsianskii (1939), in the hydrodynamic limit, the von Kármán-Howarth equation admits the invariant
$\Lambda^{{\bf u}{\bf u}}=\int_{0}^{\infty}\textrm{d}r\;r^{4}C_{r\;r}^{{\bf u}{\bf u}}(r,t)\;.$ (80)
As shown by Chandrasekhar Chandrasekhar (1951), this quantity should also be conserved in MHD turbulence, which can be seen by multiplying Eq. (8) by $r^{4}$ and subsequent integration
$\frac{\partial}{\partial t}\int_{0}^{r}\textrm{d}r^{\prime}\;r^{\prime 4}C_{r\;r}^{\mathbf{u}\mathbf{u}}(r^{\prime},t)$
$=$ $r^{4}\left(C_{r\;r\;r}^{\mathbf{u}\mathbf{u}\mathbf{u}}(r,t)-C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)+2\nu\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{u}\mathbf{u}}(r,t)\right)\;.$ (81)
Hence, if the bracketed terms on the r.h.s. decay more rapidly than $\sim r^{-4}$ in the limit of $r\rightarrow\infty$, then $\Lambda^{{\bf u}{\bf u}}$ is a conserved quantity in MHD turbulence as well. Nonetheless, due to the anti-symmetry of the third-order correlation in Eq. (9), one cannot establish a similar line of reasoning for the magnetic Loitsiansky invariant
$\Lambda^{{\bf h}{\bf h}}=\int_{0}^{\infty}\textrm{d}r\;r^{4}C_{r\;r}^{{\bf h}{\bf h}}(r,t)\;.$ (82)
Following Chandrasekhar Chandrasekhar (1951), we introduce the vector potential ${\bf a}({\bf x},t)$ according to
${\bf b}({\bf x},t)=\nabla\times{\bf a}({\bf x},t)\;,$ (83)
The anti-symmetric tensor (64) can thus be re-cast as
$A_{[ki]j}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({\bf r},t)=\langle(u_{k}({\bf x},t)h_{i}({\bf x},t)-u_{i}({\bf x},t)h_{k}({\bf x},t))h_{j}({\bf x}^{\prime},t)\rangle$
$=$ $\varepsilon_{jlm}\frac{\partial}{\partial r_{l}}\langle(u_{k}({\bf x},t)h_{i}({\bf x},t)-u_{i}({\bf x},t)h_{k}({\bf x},t))a_{j}({\bf x}^{\prime},t)\rangle\;.$ (84)
The corresponding tensor is skew and possesses the following form
$A_{[ki]j}^{\mathbf{u}\mathbf{h}{\mathbf{a}}}({\bf r},t)=\langle(u_{k}({\bf x},t)h_{i}({\bf x},t)-u_{i}({\bf x},t)h_{k}({\bf x},t))a_{j}({\bf x}^{\prime},t)\rangle$
$=$ $2A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{a}}}({r},t)\varepsilon_{ijk}+r\frac{\partial A_{[rt],t}^{\mathbf{u}\mathbf{h}{\mathbf{a}}}({r},t)}{\partial r}\left(\frac{r_{i}}{r}\varepsilon_{jkl}\frac{r_{l}}{r}+\frac{r_{k}}{r}\varepsilon_{jil}\frac{r_{l}}{r}\right)\;.$ (85)
Inserting this tensorial form in Eq. (84) thus yields
$A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({r},t)=-\frac{1}{r^{3}}\frac{\partial}{\partial r}\left(r^{4}\frac{\partial A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{a}}}({r},t)}{\partial r}\right)\;,$ (86)
which upon insertion into Eq. (9) yields
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)=\frac{1}{r^{4}}\frac{\partial}{\partial r}\left[r^{4}\left(\frac{\partial A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{a}}}({r},t)}{\partial r}+2\lambda\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\right)\right]\;.$ (87)
Hence, similar arguments as the ones used in Eq. (81) lead to $\lambda^{{\bf h}{\bf h}}=const.$, and Chandrasekhar further concluded $\lambda^{{\bf h}{\bf h}}=0$. From Eq. (9), we can thus derive
$\int_{0}^{\infty}\textrm{d}r^{\prime}r^{\prime 3}A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({r},t)=0\;.$ (88)
Appendix F Longitudinal and Transverse Correlation Functions
In this Appendix, we derive tensorial forms for second and third-order correlation functions under the assumptions of homogeneity and isotropy. To this end, we consider the velocity fields ${\bf u}({\bf x}+{\bf r},t)$ at point ${\bf x}+{\bf r}$ and ${\bf u}({\bf x},t)$ at point ${\bf x}$ (a similar treatment applies for magnetic field). We can divide the vector ${\bf u}={\bf u}_{r}+{\bf u}_{t}$ into a part ${\bf u}_{r}$ parallel to ${\bf r}$, and a transverse part ${\bf u}_{t}$. These parts are thereby given as
${\bf u}_{r}=\frac{{\bf r}}{r}\left(\frac{{\bf r}}{r}\cdot{\bf u}\right)\;,\;\;\textrm{and}\;\;{\bf u}_{t}=-\left(\frac{{\bf r}}{r}\times\left(\frac{{\bf r}}{r}\times{\bf u}\right)\right).$ (89)
The longitudinal correlation function
$C_{r\,r}(r,t)=\langle{\bf u}_{r}({\bf x},t)\cdot{\bf u}_{r}({\bf x}+{\bf r},t)\rangle\;,$ (90)
can be calculated by multiplying the two-point correlation tensor
$C_{i\,j}({\bf r},t)=\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\rangle\;,$ (91)
by $\frac{r_{i}}{r_{j}}$ and $\frac{r_{j}}{r}$. Assuming that $C_{i\,j}({\bf r},t)=\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\rangle$ is isotropic and mirror-symmetric Monin and Yaglom (2007) its general form is given by
$C_{i\,j}({\bf r},t)=C_{1}(r,t)\frac{r_{i}r_{j}}{r^{2}}+C_{2}(r,t)\delta_{ij}\;,$ (92)
where the defining scalars $C_{1}(r,t)$ and $C_{2}(r,t)$ can now be expressed in terms of the longitudinal (90) and transverse correlation functions
$C_{t\,t}(r,t)=\frac{1}{2}\langle{\bf u}_{t}({\bf x}+{\bf r},t)\cdot{\bf u}_{t}({\bf x},t)\rangle\;.$ (93)
Here, the factor $\frac{1}{2}$ has been included in three dimensions: two transverse and only one longitudinal direction. Multiplying Eq. (92) by $\frac{r_{i}r_{j}}{r^{2}}$ thus yields
$C_{rr}(r,t)=C_{1}(r,t)+C_{2}(r,t)\;.$ (94)
Decomposing the correlation tensor (98) yields
$C_{ij}({\bf r},t)=$ $\langle{\bf u}_{r}({\bf x},t)\cdot{\bf u}_{r}({\bf x}+{\bf r},t)\rangle\frac{r_{i}r_{j}}{r^{2}}$
$+\langle{u}_{t,i}({\bf x}+{\bf r},t){u}_{t,j}({\bf x},t)\rangle\;.$ (95)
Hence, we can identify
[…]
$=$ $C_{2}(r,t)\left(\delta_{ij}-\frac{r_{i}r_{j}}{r^{2}}\right)\;.$ (96)
Summing Eq. (96) over $i=j$ thus yields
$2C_{tt}(r,tt)=2C_{2}(r,t)\;,$ (97)
and we can express the correlation tensor (92) in terms of its longitudinal and transverse correlation functions according to
$C_{i\,j}({\bf r},t)=\left(C_{r\,r}(r,t)-C_{t\,t}(r,t)\right)\frac{r_{i}r_{j}}{r^{2}}+C_{t\,t}(r,t)\delta_{ij}.$ (98)
The bilinear form (92) can now be extended to a trilinear form Robertson (1940). A general isotropic and mirror-symmetric tensor of order three can be defined as
$C_{ijk}({\bf r},t)=$ $C_{1}(r,t)\frac{r_{i}r_{j}r_{k}}{r^{3}}+C_{2}(r,t)\frac{r_{k}}{r}\delta_{ij}$
$+C_{3}(r,t)\frac{r_{j}}{r}\delta_{ik}+C_{4}(r,t)\frac{r_{i}}{r}\delta_{jk}\;.$ (99)
Whereas a bilinear form is always symmetric in $i$ and $j$, we can now impose further symmetry conditions on this tensor. E.g., the third-order correlation function
$C_{(ij)k}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)=\langle u_{i}({\bf x},t)u_{j}({\bf x},t)u_{k}({\bf x}+{\bf r},t)\rangle\;,$ (100)
is symmetric in $i$ and $j$, which implies $C_{3}(r,t)=C_{4}(r,t)$. This also applies to the third-order correlation $C_{(ij)k}^{{\bf h}{\bf h}{\bf u}}=\langle h_{i}({\bf x},t)h_{j}({\bf x},t)u_{k}({\bf x}+{\bf r},t)\rangle$, which stems from the Lorentz force. Both of these tensors can thus be expressed as
$C_{(ij)k}({\bf r},t)$ $=C_{1}(r,t)\frac{r_{i}r_{j}r_{k}}{r^{3}}+C_{2}(r,t)\frac{r_{k}}{r}\delta_{ij}$
$+C_{3}(r,t)\left(\frac{r_{j}}{r}\delta_{ik}+\frac{r_{i}}{r}\delta_{jk}\right)\;,$ (101)
where the coefficients will be specified later on in terms of longitudinal and transverse correlation functions. In contrast, the third-order tensor
$A_{[ij]k}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}({\bf r},t)$ (102)
$=\langle(u_{i}({\bf x},t)h_{j}({\bf x},t)-u_{j}({\bf x},t)h_{i}({\bf x},t))h_{k}({\bf x}+{\bf r},t)\rangle\;,$
is anti-symmetric in $i$ and $j$, which implies $C_{1}(r,t)=C_{2}(r,t)=0$ and $C_{3}(r,t)=-C_{4}(r,t)=A_{1}(r,t)$. Hence, an anti-symmetric tensor of order three obeys the trilinear form
$A_{[ij]k}({\bf r},t)=A_{1}(r,t)\left(\frac{r_{j}}{r}\delta_{ik}-\frac{r_{i}}{r}\delta_{jk}\right)\;.$ (103)
Appendix G The Correlation Functions for Incompressible, Isotropic, and Homogeneous Fields
G.1 Correlation Functions of Second Order
Due to the incompressibility condition, it is possible to reduce the tensorial form of $C_{i\,j}({\bf r},t)$ to a dependence of the longitudinal structure function $C_{r\,r}(r,t)$ only. The incompressibility condition is used according to
$\frac{\partial}{\partial r_{i}}C_{i\,j}({\bf r},t)=\left\langle\frac{\partial u_{i}({\bf x}+{\bf r},t)}{\partial r_{i}}u_{j}({\bf x},t)\right\rangle=0\;.$ (104)
Using the relation $\frac{\partial}{\partial r_{i}}=\frac{r_{i}}{r}\frac{\partial}{\partial r}$ yields
$\frac{\partial}{\partial r_{i}}C_{i\,j}({\bf r},t)=\frac{\partial}{\partial r}(C_{r\,r}(r,t)-C_{t\,t}(r,t))\frac{r_{j}}{r}$ (105)
$+\frac{2}{r}(C_{r\,r}(r,t)-C_{t\,t}(r,t))\frac{r_{j}}{r}+\frac{\partial}{\partial r}C_{t\,t}(r,t)\frac{r_{j}}{r}=0\;,$
and results in the von Kármán-Howarth relation
$C_{t\,t}(r,t)=\frac{1}{2r}\frac{\partial}{\partial r}\left(r^{2}C_{r\,r}(r,t)\right)\;.$ (106)
The correlation function $C_{i\,j}({\bf r},t)$ can therefore be described solely in terms of the longitudinal correlation function $C_{r\,r}(r,t)$. Furthermore, summing $C_{i\,j}({\bf r},t)$ over equal indices $i=j$ we obtain
$\sum_{i=j}C_{i\,j}({\bf r},t)=$ $C_{r\,r}(r,t)+2C_{t\,t}(r,t)$
$=$ $\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{3}C_{r\,r}(r,t)\right)\;.$ (107)
G.2 Correlation Functions of Third Order
Applying the incompressibility condition
$\frac{\partial}{\partial r_{k}}C_{(ij)k}({\bf r},t)=0\;,$ (108)
to the third-order correlation function (101) yields
$\left(\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}C_{1}(r,t)\right)+2r\frac{\partial}{\partial r}\frac{C_{3}(r,t)}{r}\right)\frac{r_{i}r_{j}}{r^{2}}$
$+\left(\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}C_{2}(r,t)\right)+2\frac{C_{3}(r,t)}{r}\right)\delta_{ij}=0\;.$ (109)
Since both brackets in (109) have to vanish identically to satisfy the equation, we obtain two equations along with (101) for the three pre-factors $C_{1}(r,t)$, $C_{2}(r,t)$ and $C_{3}(r,t)$. This system of equations is solved by
$C_{1}(r,t)$ $=$ $-\frac{r^{2}}{2}\frac{\partial}{\partial r}\left(\frac{C_{r\,r\,r}(r,t)}{r}\right)\;,$ (110)
$C_{2}(r,t)$ $=$ $-\frac{C_{r\,r\,r}(r,t)}{2},$ (111)
$C_{3}(r,t)$ $=$ $\frac{1}{4r}\frac{\partial}{\partial r}\left(r^{2}C_{r\,r\,r}(r,t)\right)\;.$ (112)
Therefore, the third-order correlation function can be written in terms of $C_{r\,r\,r}(r,t)$ only
$C_{(ij)k}({\bf r},t)=-\frac{r^{2}}{2}\frac{\partial}{\partial r}\left(\frac{C_{r\,r\,r}(r,t)}{r}\right)\frac{r_{i}r_{j}r_{k}}{r^{3}}$
$+$ $\frac{1}{4r}\frac{\partial}{\partial r}\left(r^{2}C_{r\,r\,r}(r,t)\right)\left(\frac{r_{i}}{r}\delta_{jk}+\frac{r_{j}}{r}\delta_{ik}\right)-\frac{C_{r\,r\,r}(r,t)}{2}\frac{r_{k}}{r}\delta_{ij}\;.$
Appendix H Structure Functions of Incompressible, Isotropic, and Homogeneous Fields
We consider the corresponding moments of velocity increments to calculate the structure functions introduced in Sec. II.
H.1 Structure Functions of Second Order
The second-order longitudinal structure functions are defined as
$S_{r\;r}^{{\bf u}{\bf u}}(r,t)=$ $\left\langle\left(\left[{\bf u}({\bf x}+{\bf r},t)-{\bf u}({\bf x},t)\right]\cdot\frac{{\bf r}}{r}\right)^{2}\right\rangle\;,$ (114)
$S_{r\;r}^{{\bf h}{\bf h}}(r,t)=$ $\left\langle\left(\left[{\bf h}({\bf x}+{\bf r},t)-{\bf h}({\bf x},t)\right]\cdot\frac{{\bf r}}{r}\right)^{2}\right\rangle\;.$ (115)
These longitudinal structure functions can be related to the longitudinal correlation functions, which will be shown here based on the longitudinal velocity correlation function $C_{r\;r}^{{\bf u}{\bf u}}(r,t)$ defined by Eq.(90). We obtain
$\left\langle\left(\left[{\bf u}({\bf x}+{\bf r},t)-{\bf u}({\bf x},t)\right]\cdot\frac{{\bf r}}{r}\right)^{2}\right\rangle$ (116)
$=$ $\frac{r_{i}}{r}\left\langle{u}_{i}({\bf x}+{\bf r},t){u}_{j}({\bf x}+{\bf r},t)+{u}_{i}({\bf x},t){u}_{j}({\bf x},t)\right\rangle\frac{r_{j}}{r}$
$-\frac{r_{i}}{r}\left\langle{u}_{i}({\bf x}+{\bf r},t){u}_{j}({\bf x},t)-{u}_{i}({\bf x},t){u}_{j}({\bf x}+{\bf r},t)\right\rangle\frac{r_{j}}{r}.$
Due to homogeneity and isotropy, we obtain
$\left\langle u_{i}({\bf x},t)u_{j}({\bf x},t)\right\rangle=\left\langle u_{i}({\bf x}+{\bf r},t)u_{j}({\bf x}+{\bf r},t)\right\rangle=C_{i\;j}^{{\bf u}{\bf u}}({\bf 0},t)$
$\left\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\right\rangle=\left\langle u_{i}({\bf x}+{\bf r},t)u_{j}({\bf x},t)\right\rangle=C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)\;.$
Hence, we obtain
$S_{r\;r}^{{\bf u}{\bf u}}(r,t)=$ $\frac{r_{i}}{r}\left[2C_{i\;j}^{{\bf u}{\bf u}}({\bf 0},t)-2C_{i\;j}^{{\bf u}{\bf u}}({\bf r},t)\right]\frac{r_{j}}{r}$ (117)
$=$ $2\left[C_{r\;r}^{{\bf u}{\bf u}}({0},t)-C_{r\;r}^{{\bf u}{\bf u}}({r},t)\right]\;.$ (118)
The same treatment applies to the magnetic structure function of the second order.
H.2 Structure Functions of Third Order
The third-order longitudinal velocity structure function
$S^{{\bf u}{\bf u}{\bf u}}_{r\;r\;r}(r,t)=\left\langle\left(\left[{\bf u}({\bf x}+{\bf r},t)-{\bf u}({\bf x},t)\right]\cdot\frac{{\bf r}}{r}\right)^{3}\right\rangle$ (119)
can be related to the third-order longitudinal correlation function as
$=$ $\left\langle\left[{u}_{i}({\bf x}+{\bf r},t)-{u}_{i}({\bf x},t)\right]\left[{u}_{j}({\bf x}+{\bf r},t)-{u}_{j}({\bf x},t)\right]\right.$
$\left.\left[{u}_{k}({\bf x}+{\bf r},t)-{u}_{k}({\bf x},t)\right]\right\rangle\frac{r_{i}r_{j}r_{k}}{r^{3}}$
$=$ $2\left[C_{(ij)k}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)+C_{(jk)i}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)+C_{(ik)j}^{{\bf u}{\bf u}{\bf u}}({\bf r},t)\right]\frac{r_{i}r_{j}r_{k}}{r^{3}}$
$=$ $6C_{r\;r\;r}^{{\bf u}{\bf u}{\bf u}}(r,t)\;.$ (120)
Here, we used statistical homogeneity and isotropy and the relation $\left\langle u_{i}({\bf x},t)u_{j}({\bf x},t)u_{k}({\bf x},t)\right\rangle=0$.